OpenStudy (anonymous):

Find the points at which the graph of the equation has a vertical or horizontal tangent line. 4x2 + 16y2 – 72x – 96y + 404 = 0 horizontal tangents (x, y) = vertical tangents (x, y) =

OpenStudy (anonymous):

and satellite i wnated to thankyou again but my comp froze..lol..please help me on this one as well:)

OpenStudy (anonymous):

sattelite can you please help me out on this one...

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

oh man another implicit diff? hold on i will write it out on paper

OpenStudy (anonymous):

Find the points at which the graph of the equation has a vertical or horizontal tangent line. (Order your answers from smallest to largest x, then from smallest to largest y.) 4x2 + 16y2 – 72x – 96y + 404 = 0 horizontal tangents (x, y) = ( ), ( ) vertical tangents (x, y) = ( ), ( )

OpenStudy (anonymous):

oki cool...thankyou so much for this thought...a guy by the name of bahrom recomended you:)

OpenStudy (anonymous):

yeah i got it.

OpenStudy (anonymous):

horizontal tangent when x = 9

OpenStudy (anonymous):

vertical when y = 3

OpenStudy (anonymous):

put in x = 9 to find the corresponding y and vice versa

OpenStudy (anonymous):

leme try

OpenStudy (anonymous):

try the x i will try the y = 3

OpenStudy (anonymous):

k..im confused about where and how to plug it in

OpenStudy (anonymous):

i got x = 9 you need y. replace x by 9 in original equation to find y

OpenStudy (anonymous):

oki

OpenStudy (anonymous):

ugh im confused i know what you saying but im screwing up the algebra:(

OpenStudy (anonymous):

damn i get no answer. let me try again

OpenStudy (anonymous):

oki

OpenStudy (anonymous):

ok i got this \[y'=\frac{8x-72}{96-32y}\] and i am fairly certain this one is right

OpenStudy (anonymous):

because i did it twice. horizontal tangent means numerator = 0 so i set \[8x-72=0\] get \[x=9\]

OpenStudy (anonymous):

the problem is when i replace x by 9 in original equation and solve for y i get no solution

OpenStudy (anonymous):

wait but its asking for points

OpenStudy (anonymous):

ok here is the answer

OpenStudy (anonymous):

so what would the pooint sets be:S

OpenStudy (anonymous):

horizontal, x = 9, y = 1

OpenStudy (anonymous):

vertical x = 5, y = 3

OpenStudy (anonymous):

oh wait. there are two for each

OpenStudy (anonymous):

its saying thats inccorect

OpenStudy (anonymous):

maybe theres 4 sets of points

OpenStudy (anonymous):

so here is final answer: horizontal x=9, y = 1 x = 9, y = 5 vertical x = 5, y = 3 x = 13, y = 3

OpenStudy (anonymous):

it is an ellipse so two vertical and two horizontal ones algebra is a pain pain pain

OpenStudy (anonymous):

:)...yes omg thankyou

OpenStudy (anonymous):

please don't tell me that was wrong

OpenStudy (anonymous):

oh whew

OpenStudy (anonymous):

no it was right lol hence im saying thanks

OpenStudy (anonymous):

do u mind helkping me on anything els..i have a few from another section too

OpenStudy (anonymous):

i have about another 15 minutes in me